# Dirichlet process mixtures of block $g$ priors for model selection and prediction in linear models

### v1

@@ 1 Introduction
Model selection and model averaging are foundational tasks in statistics and machine learning. Associated Bayesian procedures typically rely on the computation of Bayes factors and posterior model probabilities, whose properties are heavily dependent of the choice of priors associated with the parameters of each model under consideration. This feature makes model selection and model averaging tasks difficult in situations where there is little prior information about the parameters of the model and standard “objective” or “default” priors (such as the reference prior, Berger et al., 2009) are improper. This is true even in well-studied settings, such as Gaussian linear models.
@@
The literature on noninformative priors and default Bayes factors for model selection for (generalized) linear models is extensive. Examples include $g$-priors (Zellner, 1986), mixtures of $g$-priors (Zellner & Siow, 1980; Liang et al., 2008), unit information priors (Kass & Wasserman, 1995), intrinsic Bayes factors (Berger & Pericchi, 1996), fractional Bayes factors (O’Hagan, 1995; De Santis & Spezzaferri, 2001), non-local priors (Johnson & Rossell, 2010, 2012) and power-expected-posterior priors (Fouskakis et al., 2015; Porwal & Rodríguez, 2023), among other approaches. See Forte et al. (2018) and Consonni et al. (2018) for recent reviews.
@@
Bayarri et al. (2012) describes a series of desiderata for default priors used for model selection and model averaging, with a particular focus on problems involving multiple linear regression. These include various forms of consistency and invariance, as well as predictive matching. More recently, Som (2014) and Som et al. (2016) suggested additional criteria related to the behavior of the Bayes factor as a subset of the significant coefficients grow to infinity. This setting is important because it serves as a proxy for situations in which effects sizes vary dramatically across covariates. Such situations arise often in practice, and they are arguably the kind of problem in which well-designed statistical methods can make a real difference. Som (2014) and Som et al. (2016) show that standard priors such as mixtures of $g$ priors (Liang et al., 2008) fail to satisfy their new criteria (a behavior they call the conditional Lindley paradox), and introduce mixtures of block $g$ priors that address the issue by grouping covariates into blocks and assigning different shrinkage parameters to each block of associated coefficients.
@@
One key shortcoming of the approach introduced in Som (2014) is that its performance depends critically on the fixed, predefined grouping of coefficients. Hence, in the absence of clear criteria for the a priori selection of the blocks, the methodology of Som (2014) is difficult to implement in practice. A second, perhaps less important shortcoming, is that the prior introduced in Som (2014) assumes that the blocks of coefficients are independent a priori. When there is strong colinearity between covariates associated with “large” and “small” coefficients, the independence assumption can lead to loss of efficiency. Our first contribution in this paper is to develop Dirichlet process (DP) mixtures of block $g$ priors that allow for differential shrinkage across coefficients while fully accounting for the observed correlations among predictors and treating the blocks of covariates as an unknown parameter that must be inferred from the data. Similar approaches have been suggested in the literature at least as early as in Liang et al. (2008) but, to the best of our knowledge, they have not been pursued before, perhaps because of perceived computational challenges. Our prior is inspired by a construction introduced by Finegold & Drton (2014) in the context of Gaussian graphical models, and allows for straightforward implementation using Markov chain Monte Carlo (MCMC) algorithms that require minimal ad hoc tuning. We study the properties of the prior and associated Bayes factors, and illustrate the approach using various real and simulated datasets.
@@
Because of our focus on differential shrinkage, the literature on continuous shrinkage priors is also relevant to our discussion. Examples of continuous shrinkage priors include the Student $t$-prior (Tipping, 2001), the Bayesian Lasso (Park & Casella, 2008; Hans, 2009), the Horseshoe prior (Carvalho et al., 2010), the Normal-Exponential-Gamma prior (Griffin & Brown, 2005) , the Normal-Gamma prior (Brown & Griffin, 2010), the Bayesian adaptive Lasso (Leng et al., 2014), the Dirichlet-Laplace prior (Bhattacharya et al., 2015), global-local shrinkage priors (Polson & Scott, 2012), the Beta-prime prior (Bai & Ghosh, 2018), the tail-adaptive shrinkage prior (Lee et al., 2020), the Horseshoe-pit prior (Denti et al., 2023), and the group Inverse-Gamma Gamma prior of Boss et al. (2023), among others. Continuous shrinkage priors tend to have computational advantages, can be connected to (frequentist) penalized likelihood methods, and are very effective in predictive settings. However, because they place probability zero on any one value of the parameter space, variable selection can be performed only by either looking at the coverage of posterior credible intervals or by thresholding the posterior distributions of the coefficients (e.g., see Li & Pati, 2017). Both of these procedures tend to work best in settings where enough prior information is available to establish practical significance. For this reason, the literature on continuous shrinkage priors is often considered as distinct from that on priors for model selection. A second contribution of this paper is to show that DP mixtures of $g$ priors provide a unifying framework for these two strands of the literature, with canonical methods in each of the two corresponding to special cases of our approach.
@@
The remainder of the paper is organized as follows. Section 2 introduces our notation and reviews the conditional Lindley paradox in the context of standard mixtures of $g$ priors. Section 3 introduces our proposed methodology and reviews its connections with the broader literature. In Section 4, we investigate the properties of the prior and the associated Bayes factors, with a particular emphasis on the criteria introduced in Bayarri et al. (2012) and Som et al. (2016). Section 5 discusses the computational implementation of our model. Section 6 and Section 7 illustrate the performance of our methodology in both simulated and real datasets. Finally, Section 8 discusses future directions for research.
@@ 2 Motivation: Bayesian variable selection and mixtures of $g$-priors
for an appropriate model-specific prior $f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})$. The prior $f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})$ is often factorized as
$f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})=f(\beta_{0},\sigma^{2})f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{\gamma}),$
with the parameters $(\beta_{0},\sigma^{2})$, which are shared across all models, being assigned the reference prior $f(\beta_{0},\sigma^{2})\propto\frac{1}{\sigma^{2}}$ (e.g., see Berger et al., 1998). A common choice for $f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{\gamma})$ is the so-called mixture of $g$-priors (Liang et al., 2008)
@@
where $\phi$ denotes the density of the multivariate normal distribution and $f(g\mid\boldsymbol{\gamma})$ is a suitable hyperprior on the common shrinkage parameter $g$. When this hyperprior is chosen carefully (e.g., an appropriately scaled member of the Compound Confluent Hypergeometric distribution introduced in Gordy, 1998), and under mild regularity conditions, Bayes factors based on mixtures of $g$ priors have various appealing theoretical properties (Liang et al., 2008; Bayarri et al., 2012; Li & Clyde, 2018). For example, the procedures are model selection consistent (the posterior probability of the true model converges in probability to 1 as the sample size grows) and information consistent (if the sequence of likelihood ratio tests associated with a sequence of samples of fixed size grows to infinite, then so does the Bayes Factor).
While the framework discussed above is the basis for much of the literature on Bayesian variable selection in linear models, procedures based on mixtures of $g$ priors do suffer from some undesirable properties. In particular, Som (2014) and Som et al. (2016) showed that Bayes factor based on mixtures of $g$ priors suffer from the so-called conditional Lindley paradox. Roughly speaking, this “paradox” states that, when comparing nested models, if at least one of the regression coefficients common to both models is large relative to other coefficients present only in the bigger model, the Bayes factor will place too much weight on the smaller model irrespective of the data generating model. More concretely, consider the two models,
@@
where $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ are $n\times p_{1}$ and $n\times p_{2}$ dimensional matrices such that $\boldsymbol{X}_{1}^{T}\boldsymbol{X}_{2}=\boldsymbol{0}$, $\boldsymbol{\beta}_{1}$ and $\boldsymbol{\beta}_{2}$ are $p_{1}$ and $p-p_{1}$ dimensional vectors, and $\boldsymbol{\epsilon}$ is the observational noise. Further, for fixed $n$, $p_{1}$, $p$, $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}\neq\boldsymbol{0}$ and $\boldsymbol{\epsilon}$, consider a sequence of vectors $\{\boldsymbol{\beta}_{1}(N):N\in\mathbb{N}\}$ and the associated sequence $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ such that $\boldsymbol{y}(N)=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}(N)+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$. Som et al. (2016) showed that, if $\left\|\boldsymbol{\beta}_{1}(N)\right\|\rightarrow\infty$ as $N\rightarrow\infty$, then, for the Bayes Factor $BF_{a,0}(\boldsymbol{y})$ based on the hyper-$g/n$ distribution $f(g)=(1+g/n)^{-a/2}$ (Liang et al., 2008), we have $BF_{a,0}\left(\boldsymbol{y}(N)\right)\rightarrow 0$, irrespective of $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}$ and $\epsilon$. To illustrate the paradox, we present in Figure 1 the behavior of $\log BF_{a,0}\left(\boldsymbol{y}(N)\right)$ for 100 randomly constructed triads $(\boldsymbol{X}_{1},\boldsymbol{X}_{2},\boldsymbol{\epsilon})$ where $n=100$, $p=2$, $p_{1}=1$, $\beta_{0}=0.5,\beta_{2}=1$. We see that, in every case, $\log BF_{a,0}\left(\boldsymbol{y}(N)\right)$ seems to decrease towards $\infty$- ∞.
Figure 1: Empirical illustration of the conditional Lindley paradox under hyper-$g/n$ prior. Thin grey lines correspond to 100 simulated datasets, while the thick blue line corresponds to the average.
The conditional Lindley paradox is a consequence of the use of a common shrinkage factor $g$ for all coefficients and cannot be solved through alternative choices of $f(g\mid\boldsymbol{\gamma})$. This is because, as some coefficients grow, the estimate of the common $g$ also must grow. The result is that small but non-zero coefficient end up being shrunk towards zero. To remedy this, Som (2014) propose the use of priors that allow for different shrinkage coefficients for various blocks of parameters. In the case of models $\mathcal{M}_{0}$ and $\mathcal{M}_{a}$ above, these block $g$-priors take the form
@@
where $g$, $g_{1}$ and $g_{2}$ are independent and identically distributed (i.i.d.), e.g., from a hyper-$g/n$ distribution. Som et al. (2016) showed that, under this block $g$-prior, the limit of $BF_{1,0}\left(\boldsymbol{y}(N)\right)$ as $N$ grows has a strictly positive lower bound, therefore avoiding the conditional Lindley paradox. Note, however, that implementing this strategy requires that we specify up front which groups of parameters will be assigned a common shrinkage parameter. This is problematic because the structure of the blocks can have a very big impact on the performance of the methods. Indeed, if at least one of the components of $\boldsymbol{\beta}_{2}$ goes to infinite as well, so that $0<\lim_{N\to\infty}\frac{\|\boldsymbol{\beta}_{1}\|^{2}}{\|\boldsymbol{\beta}_{2}\|^{2}}=d<\infty$ , then $\lim_{N\to\infty}BF_{a,0}(\boldsymbol{y}(N))=0$ and the Lindley paradox reappears.
Motivated by this observation, in the following section we build on these ideas to propose Dirichlet process mixtures of block $g$ priors, which allow for differential shrinkage and account for correlations across parameters while treating the block structure as a parameter to be learned from the data.
@@ 3 A new class of priors: Dirichlet process mixtures of block $g$ priors
where $\boldsymbol{\Omega}_{\boldsymbol{\gamma}}=\boldsymbol{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}$ (see Section A of the supplementary materials). The differential-shrinkage $g$ prior can be obtained by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$. Indeed, note that the standard $g$ prior is then obtained by further setting $g_{1}=g_{2}=\cdots=g_{p_{\boldsymbol{\gamma}}}=g$. Furthermore, when $g_{j}\sim\mathcal{O}(n)$ for all $j$, such differential-shrinkage $g$ prior can be considered (approximately) constant information.
A natural approach to modeling the $g_{j}$s is to assign them a parametric family that is flexible enough to encompass various tail behaviors. One example, which we adopt in the remainder of this paper, is
@@
where $\Gamma(z)=\int_{0}^{\infty}t^{z-1}\exp\left\{-t\right\}\mathop{}\!\mathrm{d}t$ denotes the standard Gamma function. Note that this family is defined for $a,b>-1$ and $\tau^{2}>0$ and corresponds to a Beta prior on $g/(\tau^{2}+g)$. It includes the hyper-$g/n$ prior (which corresponds to $-1/2\leq a\leq 0$, $b=0$ and $\tau^{2}=n$), as well as the half Cauchy distribution that underlies the Horseshoe prior (Carvalho et al., 2010) (which corresponds to $a=b=-1/2$ and assigning $\tau^{2}$ a half-Cauchy distribution). Hence, borrowing from the literature on continuous shrinkage priors, we call priors of this type “global-local” $g$ priors, where $g_{1},\ldots,g_{p_{\gamma}}$ are “local” shrinkage parameter and $\tau^{2}$ is a “global” shrinkage parameter (which can be either known or unknown).
One potential challenge of the approach just outlined is the need to estimate what is, potentially, a very large number of different shrinkage factors. This is not only computationally costly, but the data is likely to have limited information about each of them. Another challenge is that the performance of the procedure can be affected by the choice of the parametric family used to model the $g_{j}$s, which is somewhat arbitrary.
@@
where $\delta_{a}$ denotes a point mass at $a$, $g^{*}_{1},g^{*}_{2},\ldots$ is an i.i.d. sequence with $g^{*}_{k}\sim H_{0}$, and $w_{k}=v_{k}\prod_{l<k}(1-v_{l})$, with $v_{1},v_{2},\ldots$ another i.i.d. sequence with $v_{k}\sim\mathsf{beta}(1,\alpha)$ (Sethuraman, 1994). Because the samples from a Dirichlet process are almost surely discrete distributions, if $g_{1},\ldots,g_{c}$ is an i.i.d. sample from a random $H\mid H_{0},\alpha\sim\mathsf{DP}(\alpha,H_{0})$, there is a positive probability of ties among the $g_{i}$s. In fact, marginalizing over $H$, their joint distribution can be described through two sequences, $\tilde{g}_{1},\tilde{g}_{2},\ldots$ i.i.d. such that $\tilde{g}_{k}\sim H_{0}$, and $\xi_{1},\ldots,\xi_{c}$ such that $\xi_{1}=1$ and
@@
where $K^{j-1}=\max_{j^{\prime}<j}\{\xi_{j^{\prime}}\}$ and $m_{k}^{j}=\sum_{j^{\prime}<j}\mathbb{I}(\xi_{j^{\prime}}=k)$ (Blackwell & MacQueen, 1973). The value of $g_{j}$ can then be recovered from those of $\tilde{g}_{1},\tilde{g}_{2},\ldots$ and $\xi_{1},\ldots,\xi_{c}$ through the relationship $g_{j}=\tilde{g}_{\xi_{j}}$. The vector $\boldsymbol{\xi}$ defines a partition $\rho=\{S_{1},\ldots,S_{K}\}$ of the set =$\mathcal{I}=\{1,\ldots,c\}$ such that $\cup_{k=1}^{K}S_{k}=\mathcal{I}$, $S_{k}\cap S_{k^{\prime}}=\emptyset$ for $k\neq k^{\prime}$, and $|S_{k}|=m_{k}$ is the number of elements in $S_{k}$, so that $i\in S_{k}$ if and only if $\xi_{i}=k$ and
$f(\rho\mid\alpha)=\frac{\Gamma(\alpha)}{\Gamma(\alpha+c)}\alpha^{K}\prod_{k=1}^{K}\Gamma(m_{k}).$
In our setting, for $j=1,\ldots,p_{\boldsymbol{\gamma}}$, we let
$g_{j}\mid H$ $\sim H,$ $H\mid\alpha$ $\sim\mathsf{DP}(\alpha,H_{0}),$
where $H_{0}$ is set to be the distribution associated with the density in (3.3), and $\alpha$ is assigned the default prior introduced in Rodríguez (2013), which has density $f(\alpha\mid\boldsymbol{\gamma})=\sqrt{\frac{1}{\alpha}\sum_{j=1}^{p_{\boldsymbol{\gamma}}-1}\frac{j}{(\alpha+j)^{2}}}$. We call the resulting prior,
@@
a Dirichlet process mixture of block $g$ priors.

### v2

@@ 1 Introduction
Model selection and model averaging are foundational tasks in statistics and machine learning. Associated Bayesian procedures typically rely on the computation of Bayes factors and posterior model probabilities, whose properties are heavily dependent of the choice of priors associated with the parameters of each model. This feature makes model selection and model averaging tasks difficult in situations where standard “objective” or “default” priors (such as the reference prior, Berger et al., 2009) are improper. This is true even in well-studied settings, such as Gaussian linear models.
@@
The literature on noninformative priors and default Bayes factors for model selection for (generalized) linear models is extensive. Examples include $g$-priors (Zellner, 1986), mixtures of $g$-priors (Zellner and Siow, 1980; Liang et al., 2008), unit information priors (Kass and Wasserman, 1995), intrinsic Bayes factors (Berger and Pericchi, 1996), non-local priors (Johnson and Rossell, 2010, 2012) and power-expected-posterior priors (Fouskakis et al., 2015; Porwal and Rodríguez, 2023), among other approaches. See Forte et al. (2018) and Consonni et al. (2018) for recent reviews.
@@
Bayarri et al. (2012) describes a series of desiderata for default priors used for model selection and model averaging, with a particular focus on problems involving multiple linear regression. These include various forms of consistency and invariance, as well as predictive matching. More recently, Som (2014) and Som et al. (2016) suggested additional criteria related to the behavior of the Bayes factor as a subset of the significant coefficients grow to infinity. This setting is important because it serves as a proxy for situations in which effects sizes vary dramatically across covariates. Such situations arise often in practice, and they are arguably the kind of problem in which well-designed statistical methods can make a real difference. Som (2014) and Som et al. (2016) show that mixtures of $g$ priors (Liang et al., 2008) fail to satisfy their new criteria (a behavior they call the conditional Lindley paradox), and introduce mixtures of block $g$ priors, which address the issue by assigning different shrinkage parameters to preselected groups of coefficients.
@@
In the absence of prior information to drive the a priori selection of the blocks, the methodology of Som (2014) is difficult to implement in practice. Furthermore, the prior introduced in Som (2014) assumes that the blocks of coefficients are independent a priori. When there is strong colinearity between covariates associated with “large” and “small” coefficients, the independence assumption can lead to loss of efficiency. Our first contribution in this paper is to develop Dirichlet process (DP) mixtures of block $g$ priors that allow for differential shrinkage across coefficients while fully accounting for the observed correlations among predictors and treating the blocks of covariates as an unknown parameter that must be inferred from the data. Similar approaches have been suggested in the literature at least as early as in Liang et al. (2008) but, to the best of our knowledge, they have not been pursued before, perhaps because of perceived computational challenges.
@@
Because of our focus on differential shrinkage, the literature on continuous shrinkage priors is also relevant to our discussion. Examples of continuous shrinkage priors include the Student $t$-prior (Tipping, 2001), the Bayesian Lasso (Park and Casella, 2008; Hans, 2009), the Horseshoe prior (Carvalho et al., 2010), the Normal-Gamma prior (Griffin and Brown, 2005; Brown and Griffin, 2010), semiparametric multiple-shrinkage priors (MacLehose and Dunson, 2010), the Bayesian adaptive Lasso (Leng et al., 2014), the Dirichlet-Laplace prior (Bhattacharya et al., 2015), global-local shrinkage priors (Polson and Scott, 2012), the Beta-prime prior (Bai and Ghosh, 2018), the Horseshoe-pit prior (Denti et al., 2023), the group Inverse-Gamma Gamma prior of Boss et al. (2023), and global-local-tail priors (Lee et al., 2024) among others. Continuous shrinkage priors tend to have computational advantages, can be connected to penalized likelihood methods, and are very effective in predictive settings. However, because they place probability zero on any one value of the parameter space, variable selection can be performed only by either looking at the coverage of posterior credible intervals or by thresholding the posterior distributions of the coefficients (e.g., see Li and Pati, 2017). Both of these procedures tend to work best in settings where enough prior information is available to establish practical significance. For this reason, the literature on continuous shrinkage priors is often considered as distinct from that on priors for model selection. A second contribution of this paper is to show that DP mixtures of $g$ priors provide a unifying framework for these two strands of the literature, with canonical methods in each of the two corresponding to special cases of ours.
@@
The remainder of the paper is organized as follows. Section 2 introduces our notation and reviews the conditional Lindley paradox. Section 3 introduces our proposed methodology and reviews its connections with the broader literature. In Section 4, we investigate the properties of the prior and the associated Bayes factors, with a particular emphasis on the criteria introduced in Bayarri et al. (2012) and Som et al. (2016). Section 5 discusses the computational implementation of our model. Section 6 and Section 7 illustrate the performance of our methodology in both simulated and real datasets. Finally, Section 8 discusses future directions for research.
@@ 2 Motivation: Bayesian variable selection and mixtures of $g$-priors
for an appropriate model-specific prior $f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})$, which is often factorized as $f(\beta_{0},\boldsymbol{\beta}_{\boldsymbol{\gamma}},\sigma^{2}\mid\boldsymbol{\gamma})=f(\beta_{0},\sigma^{2})f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{\gamma})$. The parameters $(\beta_{0},\sigma^{2})$ are usually assigned the reference prior $f(\beta_{0},\sigma^{2})\propto\frac{1}{\sigma^{2}}$ (e.g., see Berger et al., 1998). A common choice for $f(\boldsymbol{\beta}_{\boldsymbol{\gamma}}\mid\sigma^{2},\boldsymbol{\gamma})$ is the so-called mixture of $g$-priors (Liang et al., 2008)
@@
where $\phi$ denotes the density of the multivariate normal distribution and $f(g\mid\boldsymbol{\gamma})$ is a suitable hyperprior. When $f(g\mid\boldsymbol{\gamma})$ is chosen carefully (e.g., an appropriately scaled member of the Compound Confluent Hypergeometric distribution introduced in Gordy, 1998), and under mild regularity conditions, Bayes factors based on mixtures of $g$ priors have various appealing theoretical properties like model selection consistency and information consistency (Liang et al., 2008; Bayarri et al., 2012).
In spite of these strong theoretical guarantees, procedures based on mixtures of $g$ priors do suffer from some undesirable properties. For example, Som (2014) and Som et al. (2016) showed that Bayes factor based on mixtures of $g$ priors suffer from the conditional Lindley paradox. Roughly speaking, this “paradox” states that, when comparing nested models, if at least one of the regression coefficients common to both models is large relative to other coefficients present only in the bigger model, the Bayes factor will place too much weight on the smaller model irrespective of the data generating model. Consider the two models,
@@
where $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ are $n\times p_{1}$ and $n\times p_{2}$ dimensional matrices such that $\boldsymbol{X}_{1}^{T}\boldsymbol{X}_{2}=\boldsymbol{0}$, $\boldsymbol{\beta}_{1}$ and $\boldsymbol{\beta}_{2}$ are $p_{1}$ and $p-p_{1}$ dimensional vectors, and $\boldsymbol{\epsilon}$ is the observational noise. Further, for fixed $n$, $_{1}, pp$, $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}\neq\boldsymbol{0}$ and $\boldsymbol{\epsilon}$, consider a sequence of vectors $\{\boldsymbol{\beta}_{1}(N):N\in\mathbb{N}\}$ and the associated sequence $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ such that $\boldsymbol{y}(N)=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}(N)+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$. Som et al. (2016) showed that, if $\left\|\boldsymbol{\beta}_{1}(N)\right\|\rightarrow\infty$ as $N\rightarrow\infty$, then, for the Bayes Factor for comparing $\mathcal{M}_{a}$ and $\mathcal{M}_{0}$, $BF_{a,0}(\boldsymbol{y})$, based on the hyper-$g/n$ distribution $f(g)=(1+g/n)^{-a/2}$ (Liang et al., 2008), we have $BF_{a,0}\left(\boldsymbol{y}(N)\right)\rightarrow 0$, irrespective of $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}$ and $\epsilon$. To illustrate the paradox, we present in Figure 1 the behavior of $\log BF_{a,0}\left(\boldsymbol{y}(N)\right)$ for 100 randomly constructed triads $(\boldsymbol{X}_{1},\boldsymbol{X}_{2},\boldsymbol{\epsilon})$ where $n=100$, $p=2$, $p_{1}=1$, $\beta_{0}=0.5,\beta_{2}=1$. We see that, in every case, $\log BF_{a,0}\left(\boldsymbol{y}(N)\right)$ seems to decrease towards -$\infty$.
Figure 1: Empirical illustration of the conditional Lindley paradox. Thin grey lines correspond to 100 simulated datasets, while the thick blue line corresponds to the average.
The conditional Lindley paradox is a consequence of the use of a common shrinkage factor $g$ and cannot be solved through alternative choices of the prior on a single shrinkage coefficient $g$. This is because, as some coefficients grow, the estimate of the common $g$ also must grow. The result is that small but non-zero coefficient end up being shrunk towards zero. To remedy this, Som (2014) propose priors that allow for different shrinkage coefficients for various blocks of parameters. In the case of $\mathcal{M}_{0}$ and $\mathcal{M}_{a}$ above, these are
@@
where $g$, $g_{1}$ and $g_{2}$ are independent and identically distributed (i.i.d.), e.g., from a hyper-$g/n$ distribution. Som et al. (2016) showed that, under this block $g$-prior, the limit of $BF_{1,0}\left(\boldsymbol{y}(N)\right)$ as $N$ grows has a strictly positive lower bound, therefore avoiding the conditional Lindley paradox. Note, however, that implementing this strategy requires that we specify up front which groups of parameters will be assigned a common shrinkage parameter. This is problematic because the structure of the blocks can have a very big impact on the performance of the methods. Indeed, if at least one of the components of $\boldsymbol{\beta}_{2}$ goes to infinite as well, so that $0<\lim_{N\to\infty}\frac{\|\boldsymbol{\beta}_{1}\|^{2}}{\|\boldsymbol{\beta}_{2}\|^{2}}=d<\infty$ , then $\lim_{N\to\infty}BF_{a,0}(\boldsymbol{y}(N))=0$.
@@ 3 A new class of priors: Dirichlet process mixtures of block $g$ priors
where $\boldsymbol{\Omega}_{\boldsymbol{\gamma}}=\boldsymbol{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}$ (see Section LABEL:ap:eq:marlikcond of the supplementary materials). The differential-shrinkage $g$ prior can be obtained by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$. Indeed, note that the standard $g$ prior is then obtained by further setting $g_{1}=g_{2}=\cdots=g_{p_{\boldsymbol{\gamma}}}=g$.
A natural approach to modeling the $g_{j}$s is to assign them a parametric family that is flexible enough to encompass various tail behaviors. One example is
@@
where $\Gamma(z)=\int_{0}^{\infty}t^{z-1}\exp\left\{-t\right\}\mathop{}\!\mathrm{d}t$ denotes the standard Gamma function. Note that this family is defined for a,b>1a,b>-1 and $\tau^{2}>0$ and corresponds to a Beta prior on $g/(\tau^{2}+g)$. It includes, for example, the hyper-$g/n$ prior (which corresponds to $1/2\leq a\leq 0$, $b=0$ and $\tau^{2}=n$), as well as the half Cauchy distribution that underlies the Horseshoe prior (Carvalho et al., 2010) (which corresponds to a=b=1/2a=b=-1/2 and assigning $\tau^{2}$ a half-Cauchy distribution). Hence, borrowing from the literature on continuous shrinkage priors, we call priors of this type “global-local” $g$ priors, where $g_{1},\ldots,g_{p_{\gamma}}$ are “local” shrinkage parameter and $\tau^{2}$ is a “global” shrinkage parameter (which can be either known or unknown).
One potential challenge of the approach just outlined is the need to estimate what is, potentially, a very large number of different shrinkage factors. This is not only computationally costly, but the data is likely to have limited information about each of them, especially in high-correlation settings. Another challenge is that the performance of the procedure can be affected by the choice of the parametric family used to model the $g_{j}$s. In particular, recent work by Piironen et al. (2017) and Lee et al. (2024) indicates that, in the context of continuous shrinkage priors, the optimal tail behavior of $p(g_{j}\mid\tau^{2})$ might depend on the level of sparsity among coefficients.
@@
where $\delta_{a}$ denotes a point mass at $a$, $g^{*}_{1},g^{*}_{2},\ldots$ is an i.i.d. sequence with $g^{*}_{k}\sim H_{0}$, and $w_{k}=v_{k}\prod_{l<k}(1-v_{l})$, with $v_{1},v_{2},\ldots$ another i.i.d. sequence with $v_{k}\sim\mathsf{beta}(1,\alpha)$ (Sethuraman, 1994). Because the samples from a Dirichlet process are almost surely discrete distributions, if $g_{1},\ldots,g_{c}$ is an i.i.d. sample from a random $H\mid H_{0},\alpha\sim\mathsf{DP}(\alpha,H_{0})$, there is a positive probability of ties among the $g_{i}$s. In fact, their joint distribution can be described through two sequences, $\tilde{g}_{1},\tilde{g}_{2},\ldots$ i.i.d. such that $\tilde{g}_{k}\sim H_{0}$, and $\xi_{1},\ldots,\xi_{c}$ such that $\xi_{1}=1$ and
@@
where $K^{j-1}=\max_{j^{\prime}<j}\{\xi_{j^{\prime}}\}$ and $m_{k}^{j}=\sum_{j^{\prime}<j}\mathbb{I}(\xi_{j^{\prime}}=k)$ (Blackwell and MacQueen, 1973). The value of $g_{j}$ can then be recovered from those of $\tilde{g}_{1},\tilde{g}_{2},\ldots$ and $\xi_{1},\ldots,\xi_{c}$ through the relationship $g_{j}=\tilde{g}_{\xi_{j}}$. The vector $\boldsymbol{\xi}$ defines a partition $\rho=\{S_{1},\ldots,S_{K}\}$ of the set $\mathcal{I}=\{1,\ldots,c\}$ such that $\cup_{k=1}^{K}S_{k}=\mathcal{I}$, $S_{k}\cap S_{k^{\prime}}=\emptyset$ for $k\neq k^{\prime}$, and $|S_{k}|=m_{k}$ is the number of elements in $S_{k}$, so that $i\in S_{k}$ if and only if $\xi_{i}=k$ and $f(\rho\mid\alpha)=\frac{\Gamma(\alpha)}{\Gamma(\alpha+c)}\alpha^{K}\prod_{k=1}^{K}\Gamma(m_{k})$.
In our setting, we let $g_{j}\mid H$ i.i.d. for $j=1,\ldots,p_{\boldsymbol{\gamma}}$ and $H\mid\alpha\sim\mathsf{DP}(\alpha,H_{0})$, where $H_{0}$ is the distribution associated with (3.3), and $\alpha$ is assigned the parameterization-invariant prior introduced in Rodríguez (2013), which has density $f(\alpha\mid\boldsymbol{\gamma})=\sqrt{\frac{1}{\alpha}\sum_{j=1}^{p_{\boldsymbol{\gamma}}-1}\frac{j}{(\alpha+j)^{2}}}$. The use of this prior circumvents the need to elicit prior information about the number of groups into which the coefficients are expected to be clustered. The resulting prior,
@@
is a Dirichlet process mixture of block $g$ priors.